OpenStudy (kayders1997):

Find the derivative of y=(3x^2)/(x^2+2x-1)^1/2

OpenStudy (xapproachesinfinity):

use quotient rule

zepdrix (zepdrix):

Are you on a real computer today? Can you see the math formatted correctly? \[\large\rm y=\frac{3x^2}{\sqrt{x^2+2x-1}}\]

OpenStudy (kayders1997):

I can see it

zepdrix (zepdrix):

Oo yay! \c:/

OpenStudy (kayders1997):

I know right! :)

zepdrix (zepdrix):

I usually like to "set up" the quotient rule before actually taking the derivatives. But whatever works best for you.\[\rm \left(\frac{3x^2}{\sqrt{x^2+2x-1}}\right)'=\frac{\color{royalblue}{\left[3x^2\right]'}\sqrt{x^2+2x-1}-3x^2\color{royalblue}{\left[\sqrt{x^2+2x-1}\right]'}}{\left(\sqrt{x^2+2x-1}\right)^2}\]

OpenStudy (kayders1997):

Omg...of course its not showing up

zepdrix (zepdrix):

you -_- you and your ways...

OpenStudy (kayders1997):

I see the first half

zepdrix (zepdrix):

get on a computer. get on that chrome :O

OpenStudy (kayders1997):

Okay just wait I'm at school!

OpenStudy (kayders1997):

okay I can see it all

OpenStudy (kayders1997):

so the derivative of the first part would be 6x times the square root I know that

zepdrix (zepdrix):

\[\rm =\frac{\color{orangered}{\left[6x\right]}\sqrt{x^2+2x-1}-3x^2\color{royalblue}{\left[\sqrt{x^2+2x-1}\right]'}}{\left(\sqrt{x^2+2x-1}\right)^2}\]Ok good.

OpenStudy (kayders1997):

And than idk...all I know is for the next part the x^2+2x-1 stay there because it is the chain rule because its more than just x

OpenStudy (kayders1997):

Gosh i've never been so confused with math!

zepdrix (zepdrix):

Derivative of `square root` is really really useful. You should try to put this one away in your brain somewhere, instead of converting to 1/2 power. \[\large\rm \frac{d}{dx}\sqrt x=\frac{1}{2\sqrt x}\] Derivative of square root is simply ... one over two square roots.

zepdrix (zepdrix):

But yes, we'll have some chain rule business going on, giving us some extra stuff in addition to that.

OpenStudy (kayders1997):

okay the derivative of the square root makes sense

zepdrix (zepdrix):

So we can generalize our square root derivative,\[\large\rm \frac{d}{dx}\sqrt{stuff}=\frac{1}{2\sqrt{stuff}}stuff'\]When our inner function is more than just x, we apply the chain rule like this.

zepdrix (zepdrix):

So for our problem,\[\large\rm \frac{d}{dx}\sqrt{x^2+2x-1}=\frac{1}{2\sqrt{x^2+2x-1}}(x^2+2x-1)'\]

zepdrix (zepdrix):

And what's the derivative of that last piece? :)

OpenStudy (kayders1997):

2x+2

zepdrix (zepdrix):

Ok great. Let's put it in the numerator,\[\large\rm \frac{d}{dx}\sqrt{x^2+2x-1}=\frac{(2x+2)}{2\sqrt{x^2+2x-1}}\]And then let's divide a 2 out of everything.\[\large\rm \frac{d}{dx}\sqrt{x^2+2x-1}=\frac{x+1}{\sqrt{x^2+2x-1}}\]

OpenStudy (kayders1997):

im confused lol

zepdrix (zepdrix):

With what part? :) Chain rule? Fancy fraction trick? The 2 division? All of it? :d

OpenStudy (kayders1997):

oh just kidding you took out the 2 on the top and than you cancelled the 2 and you were left with (x+1) on the top?

zepdrix (zepdrix):

Ya sorry, I should have included the intermediate step of factoring. 2x+2 = 2(x+1) and then cancelled the 2's.

OpenStudy (kayders1997):

nope thats okay, I am a blonde so it can take me a while

zepdrix (zepdrix):

\[\rm =\frac{\color{orangered}{\left[6x\right]}\sqrt{x^2+2x-1}-3x^2\color{royalblue}{\left[\sqrt{x^2+2x-1}\right]'}}{\left(\sqrt{x^2+2x-1}\right)^2}\]Mmm k good, that takes care of our other blue piece for us,\[\rm =\frac{\color{orangered}{\left[6x\right]}\sqrt{x^2+2x-1}-3x^2\color{orangered}{\left[\dfrac{x+1}{\sqrt{x^2+2x-1}}\right]}}{\left(\sqrt{x^2+2x-1}\right)^2}\]

zepdrix (zepdrix):

Looks like we can do something with the denominator, ya? Do you see how that will simplify?

OpenStudy (kayders1997):

it will be just what it is without the square root

OpenStudy (kayders1997):

wait hold on a minute I got rid of the -3x^2 for some reason?

zepdrix (zepdrix):

Yes, square root and square undo one another :)\[\rm =\frac{6x\sqrt{x^2+2x-1}\quad-\quad\dfrac{3x^2(x+1)}{\sqrt{x^2+2x-1}}}{x^2+2x-1}\]

zepdrix (zepdrix):

you did? Hmm that's no bueno :O Try to backtrack and see what happened to it.

OpenStudy (kayders1997):

I still had it in there I just lost it for some odd reason, and lol love the espanol

zepdrix (zepdrix):

lol :D Ok so now... let's clean it up a tad bit. We'll get rid of the fraction on fraction business by multiplying through by the "upper" denominator.\[\rm =\frac{6x\sqrt{x^2+2x-1}\quad-\quad\dfrac{3x^2(x+1)}{\sqrt{x^2+2x-1}}}{x^2+2x-1}\left(\frac{\sqrt{x^2+2x-1}}{\sqrt{x^2+2x-1}}\right)\]

OpenStudy (kayders1997):

And than the square roots can go away?

OpenStudy (kayders1997):

wait what?

zepdrix (zepdrix):

The one in the denominator? Yes. That's the point of applying this trick. Lemme add some color, just to make sure it's super clear.

zepdrix (zepdrix):

\[\rm =\frac{6x\sqrt{x^2+2x-1}\quad-\quad\dfrac{3x^2(x+1)}{\sqrt{x^2+2x-1}}}{x^2+2x-1}\left(\frac{\color{orangered}{\sqrt{x^2+2x-1}}}{\sqrt{x^2+2x-1}}\right)\]This orange one will be distributed to each term in the numerator. The other one will simply go to the bottom.\[\rm =\frac{6x\sqrt{x^2+2x-1}\color{orangered}{\sqrt{x^2+2x-1}}\quad-\quad\color{orangered}{\sqrt{x^2+2x-1}}\dfrac{3x^2(x+1)}{\sqrt{x^2+2x-1}}}{(x^2+2x-1)\sqrt{x^2+2x-1}}\]

zepdrix (zepdrix):

Uh oh, I smell brain. Brain esplode? I hate when that happens :(

OpenStudy (kayders1997):

I'm so confused

OpenStudy (kayders1997):

so...? Do the square roots cancel in the second term?

zepdrix (zepdrix):

Yes, they cancel :) That's why we're doing this trick. To get rid of that nasty fraction.

OpenStudy (kayders1997):

okay that makes sence now

zepdrix (zepdrix):

\[\large\rm =\frac{6x\sqrt{x^2+2x-1}\color{orangered}{\sqrt{x^2+2x-1}}\quad-\quad3x^2(x+1)}{(x^2+2x-1)\sqrt{x^2+2x-1}}\]So yah that whole mess cleans up really really nicely in the second term.

OpenStudy (kayders1997):

and than the square roots cancel again? hmmm? Am I understanding this right?

OpenStudy (kayders1997):

ohhhhhhh the square root is multiplied by both terms?

zepdrix (zepdrix):

In the numerator? Each `term` got one of the orange things. The one with the 3x^2(x+1) cancelled out. So now only the first term has one.

zepdrix (zepdrix):

And ya, the square roots will go away, sort of :) We'll have to do something clever in the denominator though.

zepdrix (zepdrix):

\[\large\rm \sqrt{stuff}\cdot\sqrt{stuff}=stuff\]It just sort of goes away in the numerator though, ya?

OpenStudy (kayders1997):

Ohhhhhhhhh, this is to much work

OpenStudy (kayders1997):

so your stuck with 6x(x^2+2x-1)-3x^2(x+1) on the top?

zepdrix (zepdrix):

\[\large\rm =\frac{6x(x^2+2x-1)\quad-\quad3x^2(x+1)}{(x^2+2x-1)\sqrt{x^2+2x-1}}\]Mmm looks good. How bout in the bottom? It actually benefits us to write this root as a 1/2 power at this point,\[\large\rm =\frac{6x(x^2+2x-1)\quad-\quad3x^2(x+1)}{(x^2+2x-1)^1(x^2+2x-1)^{1/2}}\]

zepdrix (zepdrix):

Do you remember how to simplify something like this?\[\large\rm x^1\cdot x^{1/2}\]It requires your exponent rules.

OpenStudy (kayders1997):

so distribute on the top? and than for the bottom don't you have to add the exponents?

OpenStudy (kayders1997):

I'm gonna have to go for a bit I'll be back in about 10-15 mins

zepdrix (zepdrix):

Good good good :) You add the exponents.\[\large\rm =\frac{6x(x^2+2x-1)\quad-\quad3x^2(x+1)}{(x^2+2x-1)^{3/2}}\]

zepdrix (zepdrix):

And then yes, distribute on the top, combine like-terms. Your teacher is really mean by the way, giving you problems like this. It's just... wrong. not cool prof.

OpenStudy (kayders1997):

I know right, uff I glad I made it home safe, and it is really hard